OpenStudy (anonymous):

find m(x) if m(0) = 8 and m'(x) = x^2 * e^(x^3)

OpenStudy (anonymous):

a u sub should get it put \(u=x^3, du=3x^2dx\) and you are almost done

OpenStudy (anonymous):

@satellite73 help me with my question next please?

OpenStudy (anonymous):

you get \[\frac{1}{3}\int e^udu = \frac{1}{3}e^u+C=\frac{1}{3}e^{x^3}+C\] then solve for \(C\) since \[m(0)=\frac{1}{3}+C=8\]

OpenStudy (anonymous):

@brittneyy18rm i'll look at it

OpenStudy (anonymous):

Thanks a bunch!